4x – 5 = 7 does not have an integer as its solution.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

The Given statement isFalse
Consider the equation, 4x–5 = 7
Transpose -5 to RHS then it becomes 5.
4x = 7 + 5 4x = 12
x = 12/4
x= 3

Thanks